# A Multimodal Lightweight Approach to Fault Diagnosis of Induction Motors in High-Dimensional Dataset

### v1

@@ Abstract
Index Terms:
induction motor, transfer learning, lightweight deep learning models, fault diagnosis, Spectral Images
@@ I Introduction
In recent years, multiple methodologies have been employed for detecting and categorizing bearing, stator, and BRB faults in IMs. These encompass thermal, induced voltage, variation of torque, vibration, and motor current signature analysis (MCSA)[8],[9]. Similarly, various ML and DL methods have been used to identify and classify these faults in IMs. These included logistic regression (LR), k-nearest Neighbors (KNN), support vector machines (SVM), decision trees (DT), ensemble learning, artificial neural networks (ANN), convolution neural networks (CNN), and recurrent neural networks (RNN) [1].
@@ II-A Fast Fourier Transform (FFT)
The FFT is a computational algorithm specifically crafted for computing the Discrete Fourier Transform (DFT) of a given signal. Fourier analysis involves the conversion of a signal from its original domain, usually time, into a representation within the frequency domain. The DFT is obtained by decomposing a sequence of values into constituent components characterized by different frequencies. If we have a sequence of complex numbers represented as $x_{0},x_{1},x_{2},\ldots,x_{n-1}$, each $x_{i}$ denotes an element in the sequence, where $i$ ranges from 0 to $n-1$. The formula defining DFT is expressed in equation 1:
@@ II-B Short-Time Fourier Transform (STFT)
Where, $X[m,i]$ is the STFT coefficient at index $m$ and frequency index $i$, $x[z]$ is the input signal, $w[z-mR]$ is the window function centered at time index $m$ with a hop size of $R$ between successive windows, $Z$ is the length of the DFT for each window.
@@ II-D Evaluation Measures
[…]
@@ V Results and Discussion
Model Name Motor Condition Classification Report
@@
Actual Labels BRB1 1850 20 13 117 0 Actual Labels BRB1 1822 10 21 147 0
@@
Actual Labels BRB1 1735 14 4 247 0 Actual Labels BRB1 1865 9 12 114 0
@@
Actual Labels BRB1 1869 8 17 106 0 Actual Labels BRB1 1909 2 6 83 0

### v2

@@ I Introduction
In recent years, multiple methodologies have been employed for detecting and categorizing bearing, stator, and BRB faults in IMs. These encompass thermal, induced voltage, variation of torque, vibration, and motor current signature analysis (MCSA)[8],[9]. Similarly, various ML [10] and DL[11] methods have been used to identify and classify these faults in IMs. These included logistic regression (LR), k-nearest Neighbors (KNN), support vector machines (SVM), decision trees (DT), ensemble learning[11], artificial neural networks (ANN)[12], convolution neural networks (CNN), and recurrent neural networks (RNN) [1].
@@ II-A Fast Fourier Transform (FFT)
The FFT is a computational algorithm specifically crafted for computing the Discrete Fourier Transform (DFT) of a given signal. Fourier analysis involves the conversion of a signal from its original domain, usually time, into a representation within the frequency domain. The DFT is obtained by decomposing a sequence of values into constituent components characterized by different frequencies. If we have a sequence of complex numbers represented as $x_{0},x_{1},x_{2},\ldots,x_{n-1}$, each $x_{i}$ denotes an element in the sequence, where $i$ ranges from 0 to n1n-1. The formula defining DFT is expressed in equation 1:
@@ II-B Short-Time Fourier Transform (STFT)
Where, $X[m,i]$ is the STFT coefficient at index $m$ and frequency index $i$, $x[z]$ is the input signal, w[zmR]w[z-mR] is the window function centered at time index $m$ with a hop size of $R$ between successive windows, $Z$ is the length of the DFT for each window.
@@ II-D Evaluation Measures
$\text{F-1 score}=\frac{2*(\text{Pr}\times\text{Rc})}{\text{Pr}+\text{Rc}}$ (11)
@@ V Results and Discussion
Model Name
Motor
Condition
Classification Report
@@
Actual
Labels
BRB1 1850 20 13 117 0
Actual
Labels
BRB1 1822 10 21 147 0
@@
Actual
Labels
BRB1 1735 14 4 247 0
Actual
Labels
BRB1 1865 9 12 114 0
@@
Actual
Labels
BRB1 1869 8 17 106 0
Actual
Labels
BRB1 1909 2 6 83 0